# TF-Mamba: A Time-Frequency Network for Sound Source Localization

### v1

@@ VI Conclusion
In this paper, we introduced TF-Mamba, a novel architecture for SSL that leverages the Mamba SSM to fuse time and frequency features effectively. By replacing traditional LSTM blocks with bidirectional Mamba, TF-Mamba enhances the modeling of audio sequence contexts while maintaining computational efficiency. Our experimental results, conducted on both simulated and real-world datasets, demonstrate that TF-Mamba significantly outperforms existing state-of-the-art methods. These findings highlight the potential of SSM like Mamba in advancing SSL tasks, marking the first application of such models in this field.
VII Acknowledgement
The authors would like to thank Yabo Wang et. al. (The authors of FN-SSL [18]) for their valuable suggestions during this work on reproducing the FN-SSL model and other baseline models.

### v2

@@ 6 Conclusion
In this paper, we introduced TF-Mamba, a novel architecture for SSL that leverages the Mamba to fuse time and frequency features effectively. By replacing traditional LSTM blocks with bidirectional Mamba, TF-Mamba enhances the modeling of audio sequence contexts while maintaining computational efficiency. Our experimental results, conducted on both simulated and real-world datasets, demonstrate that TF-Mamba significantly outperforms existing state-of-the-art methods. These findings highlight the proposed TF-Mamba in advancing SSL tasks, marking the first application of such models in this field.